Assume a 10-year, 7%, commercial mortgage loan with a $1,400,000 remaining balance is paid off in 3 years. Assume interest rates have declined to 5% since origination. If the $1,400,000 is invested for the remaining 7 years at 5%, to what amount will the principal balance increase if interest is compounded monthly? What if the lump-sum is invested at a 7% annual rate with monthly compounding?
Answer #1

1. If $1,400,000 principal balance is reinvested at 5% compounded monthly, for the remaining 7 years, balance will increase to $1,985,804 (N= 7*12; i= 5/12; PV= -1,400,000; PMT=0 FV= 1,985,804)

2. If compounded monthly at 7%, principal balance = $1,400,000 x [1 + (0.07/12)](7*12)

= $1,400,000 x 1.63 = $2,281,992